OpenStudy (anonymous):

Find the equation of the planes parallel to the plane x-2y+2z-3=0 whose perpendicular distance from (1,2,3) is 1.

OpenStudy (rsadhvika):

any plane parallel to \(x-2y+2z-3 = 0\) willl be of form \(x-2y+2x+k=0\)

OpenStudy (rsadhvika):

since you're given the \(\perp\) distance from point (1, 2, 3), there can be two planes i think you can find them

OpenStudy (anonymous):

how can i find?

OpenStudy (rsadhvika):

\(\perp \) distance from a point \((x_1, y_1, z_1)\) to a plane \(ax+by+cz+d = 0\) is \(\large \frac{ax_1 + by_1+cz_1 + d}{\sqrt{a^2+b^2+c^2}}\)

OpenStudy (rsadhvika):

set it to 1 and find k

OpenStudy (anonymous):

ok and then i have to put the value of k in dat eqn right?

OpenStudy (anonymous):

ok m wrking i tout

OpenStudy (rsadhvika):

yes

OpenStudy (anonymous):

thank u! :)

OpenStudy (rsadhvika):

np :) think a bit more.... im not sure whether there will be only 1 plane or 2 planes

OpenStudy (anonymous):

hey m getting k=root14-3

OpenStudy (anonymous):